### v4

@@ 5.3.2 Allocations and Past Relationship
Next, we test the Hypothesis 2-3. A sponsor builds a relationship with non-premium investors by providing MLOT from good deals, so if and when the new target firm is weak, it expects the investors with high past MLOT, i.e., the relationship investors, to return the favor by investing in the current deal. To this end, we evaluate if a sponsor is more likely to use non-premium investors if their past MLOT from the sponsor is high and the current target is weak.
@@
Note: Columns (1) to (4) of this table show the MLE of participation (5), and column (5) shows the estimates of the allocation (6). Each observation is at (SPAC, sponsor, non-premium PIPE investor) level. For each SPAC, non-premium investors include all relationship investors who participate in at least one SPAC deal initialized by the sponsor during the sample period. For the first four columns, the dependent variable, , equals one if the non-premium investor participates in the focal SPAC deal. The dependent variable for the last column, , is the investor’s allocation as a fraction of all relationship investors’ allocations. The key independent variables include the non-premium investor’s MLOT earned with this sponsor in the past in log form, , and a measure of the (reciprocal) of the deal quality, Cold. For the allocation regression, we also include the estimated probability of participation using estimates from column (4) as an additional variable. All continuous variables at winsorized at percentile. Robust standard errors are clustered at the sponsor-PIPE pair level and reported in parentheses. For column (5), Bootstrap standard errors (500 replications) are reported in the square brackets. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
One possible concern with the estimation is that it ignores the outside option of the sponsor, which may be correlated with the unobserved error. A better outside option lowers the probability of issuing securities to a non-premium investor and vice versa. Ignoring outside options could introduce attenuation bias if cold deals have profitable outside options. To alleviate this concern, we use the highest MLOT among all premium PIPE investors not included in the current deal as an additional control variable to capture the strength of the outside option. We also include the log of IPO proceeds, days remaining, and the number of premium investors as additional covariates.
@@
Next, we examine if, conditional on participation, the sponsor allocates a higher portion of securities to non-premium investors with a stronger relationship when a deal is cold and vice versa. To this end, we determine the allocation of SPAC ’s securities from sponsor to PIPE investor among all relationship investors , , as the ratio of ’s shares to the total shares allocated to . We set the share to zero if a non-premium investor does not participate in a deal. The mean allocation is , and the standard deviation is . Furthermore, the allocation is skewed towards zero, with the median of and the -percentile at . Given that there are 1,792 non-premium investors, and so the allocation in our sample is skewed to the left with a mass at zero, we estimate the following Tobit model:
@@ 5.3.3 Relationship Capital and Liquidation Risk
We have shown that the relationship investors support the sponsor by participating in a weak deal and purchasing a larger fraction of shares. These non-premium investors received larger MLOTs from the same sponsor in the previous deals. Therefore, it stands to reason that a sponsor builds these relationships with an eye toward weak deals in the future. A sponsor with more relationship capital can afford to take more risk by working with a weak target firm because it can draw down the relationship capital.
@@
Next, we test Hypothesis 2-4 by estimating the liquidation probability as a function of MLOT earned by premium and non-premium investors in previous deals initiated by the same sponsor using a linear probability model. In our sample, we observe 191 liquidations and 473 successful mergers, and we want to test if the probability of liquidation decreases with the sponsor’s “relationship payments” to non-premium investors.
@@
We regress the binary variable that is equal to one if the SPAC is liquidated and zero otherwise, on the past MLOT, the premium and non-premium investors have made from the current sponsor and several other controls.
| (1) | (2) | (3) | (4) | (5) | |
| Premium Investors | 0.057** | 0.049 | 0.054* | 0.059* | 0.060* |
| (0.029) | (0.033) | (0.032) | (0.033) | (0.033) | |
| Non-premium Investors | -0.064** | -0.075** | -0.077** | -0.079*** | -0.081*** |
| (0.029) | (0.030) | (0.030) | (0.030) | (0.031) | |
| Controls | ✓ | ✓ | ✓ | ✓ | ✓ |
| IPO Month FEs | ✗ | ✓ | ✓ | ✓ | ✓ |
| Target Sector FEs | ✗ | ✗ | ✓ | ✓ | ✓ |
| Jurisdiction FEs | ✗ | ✗ | ✗ | ✓ | ✓ |
| Sponsor Quality | ✗ | ✗ | ✗ | ✗ | ✓ |
| Adj. | 0.047 | 0.121 | 0.127 | 0.126 | 0.123 |
| Obs. | 806 | 781 | 758 | 757 | 757 |
Note: This table demonstrates that SPACs initiated by sponsors with a history of stronger relationships with non-premium investors are less likely to liquidate when a target is announced. All regressions use linear probability models. Dependent variable is a dummy variable that is one if the SPAC liquidates and zero otherwise. The primary independent variables are the total MLOT earned by premium and non-premium investors in previous SPACs initiated by the same sponsor. Controls include (log) IPO proceeds, number of past successful deals, dummy variable if there were PIPE (pre) investors, and number of active SPACs. Sponsor quality is the sponsor’s fixed effect estimated from (2). All continuous variables are winsorized at the and percentiles. Robust standard errors, clustered at the sponsor level, are reported in parentheses. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
@@
The results from this estimation exercise are in Table 11. Column (1) suggests that SPACs initialized by sponsors who have built a stronger relationship with non-premium investors are less likely to liquidate. We progressively introduce additional control variables from columns (2) to (5). In column (2), we include SPAC’s IPO month fixed effect to capture time differences. In column (3), we include fixed effects for the SPAC’s targeted sector at its IPO to capture systematic differences across sectors that may affect the liquidation risk. In column (4), we incorporate SPAC’s jurisdiction fixed effects based on the jurisdiction the SPAC is registered in to capture differences in the legal systems that may influence the liquidation rules. Finally, in column (5), we included the sponsor fixed effect we estimated from Equation (2) as a proxy for the sponsor’s quality.
Column (5), our preferred specification, suggests that doubling the sponsor’s relationship payments to non-premium investors decreases liquidation probability by 8.1%, which is 33.6% of the sample average liquidation rate, and this estimated effect is statistically significant at the 1% confidence level.141414 Although not reported, we verify that even when we use the competing-risk model to account for right censoring because 16% of the announced deals in our sample are still searching for a target, our finding holds, albeit the estimates are slightly nosier than in Table 11.
@@
In summary, we find evidence consistent with these conditions: (i) premium PIPE investors generate value-relevant information, (ii) non-premium PIPE investors are paid agency fees, and (iii) the non-premium investors help close weaker deals in the future. In Appendix A, we show that these results are robust to removing outlier investors, placebo treatments, and the definition of premium PIPE investors.
6 Observed and Unobserved Heterogeneity
In this section, we determine how much of the difference in average earnings between premium and non-premium investors can be explained by differences in their observable characteristics. If they have different “business models” (information production vs. agency service), we expect observed heterogeneity to explain a smaller portion of the difference.
To minimize the effect of outliers in this exercise, we exclude seven out of 473 successful SPACs with closing share prices higher than $100.151515These SPACs are: Queen’s Gambit Growth Capital, Insurance Acquisition Corp., Matlin & Partners Acquisition Corporation, Northern Genesis Acquisition Corp. II, VPC Impact Acquisition Holdings III, Inc., Reinvent Technology Partners Z (f/k/a Reinvent Technology Partners B), and Ajax I. In Figure 2 (left panel), we display the histogram of total MLOT paid to all PIPE (at) investors across SPACs that have successfully merged with a private target firm. The average MLOT per deal is $62.38 million, with a large standard deviation of $199.07 million. These earnings are divided among, on average, 10.76 PIPE investors (see Table 2), of which 3.02 are premium investors. In Figure 2 (right panel), we show the shares of premium PIPE investors (left vertical line) and the total MLOT (right vertical line). Averaging these shares, we find that premium PIPE investors get (around) 26% of the MLOT. In other words, 26% of the underpricing at the time of the business closing announcement is for information production, and 74% is for quid pro quo.
Note: The left panel is the histogram of total MLOT for all investors. The right panel shows the premium PIPE investors’ share of MLOT (vertical axis on the left) and the total MLOT in that deal (vertical axis on the right).
Considering each investor separately, the MLOT of a premium investor, on average, is $6.96 million, and that of a non-premium investor is $3.17 million. Next, we determine what part of this $3.79 million difference in average MLOT can be explained by differences in the observable characteristics of these two types of investors.
First, we introduce new variables capturing some of the observed differences between these two groups. In particular, for each PIPE investor and each SPAC, we determine the number of SPACs they participated in before the current one, how much they invested, the number of past SPACs that involved a sponsor different from the current one, total MLOT from other sponsors, the number of SPACs with sponsors different from the current one, and whether the target is in the same industry as the current target. These variables capture market-wide experience and successes.
We have 4,152 deals at the SPAC and PIPE (at) investor level, of which 1,245 are with premium investors and the rest with non-premium investors. Table 12 presents the descriptive statistics and the results from the tests of mean differences between premium and non-premium investors.
Premium PIPE Sample |
Non-Premium PIPE Sample |
Difference |
MLOT ($mm) |
6.95 |
3.17 |
3.78 |
| SPAC Characteristics | |||
IPO Proceeds ($mm) |
397.25 |
363.87 |
33.38 |
Days Remaining |
492.98 |
467.59 |
25.38 |
| PIPE Investor Characteristics | |||
No. of past SPACs |
23.99 |
5.51 |
18.48 |
Past Investment Amount ($mm) |
460.09 |
50.77 |
409.32 |
No. of SPACs with other sponsors |
24.40 |
5.44 |
18.95 |
Past MLOT with other sponsors ($mm) |
292.85 |
24.61 |
268.25 |
No. of SPACs with other sponsors, same industry |
3.59 |
0.80 |
2.79 |
Past MLOT with other sponsors, same industry ($mm) |
45.83 |
3.65 |
42.18 |
Note: This table reports means and differences in our sample’s average characteristics for premium and non-premium investors. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
The two groups display distinct MLOTs and characteristics. Specifically, premium investors are involved in larger SPACs requiring fewer days to identify a target. Moreover, these investors possess greater experience in the SPAC market, as evidenced by a higher number of previous SPAC participations (23.99 vs. 5.51) and a larger investment amount ($460.09 million vs. $50.77 million). They also work with more SPACs (24.40 vs. 5.44) and obtain a higher MLOT ($292.85 million vs. $24.61 million) from other sponsors. In addition, premium investors operate more in the same industry with different sponsors (3.59 vs. 0.80).
Next, we use the Kitagawa-Blinder-Oaxaca decomposition to document the extent and drivers of the MLOT differences between premium and non-premium investors. In particular, we begin with the following: where is the vector of observed characteristics in Table 12, and is a mean zero random variable. Let , where denotes the sample average of the group . We can then re-write as
The explained portion is the portion of the MLOT difference explained by differences in observed characteristics between premium and non-premium investors. In other words, if we change to , it measures how much additional MLOT a non-premium investor gets. The unexplained portion is the part driven by differences in the coefficients.
Note: The figure displays the explained components of a twofold Kitagawa-Blinder-Oaxaca decomposition of the premium vs. non-premium MLOT. The x-axis is the mean difference in MLOT (in millions), and the y-axis lists observable characteristics.
We find that the mean increase in non-premium investors’ MLOT is $0.04 million (s.e. $1.38 million) if they had the same characteristics as premium investors. So, the differences in observed characteristics do not account for the MLOT gap. Figure 3 displays how much each of the eight components in explains, along with their confidence intervals. We find that the difference in average MLOT cannot be explained by differences in average , suggesting these two types of investors have different business models.
@@ 7 Conclusion
Understanding reasons for underpricing hinges on the relative importance of information production and quid pro quo. In this paper, we determine the relative importance within the context of SPACs. Distinguishing between premium and non-premium PIPE (institutional) investors, we show that the former produce information and, on average, are paid 26% of the total underpricing. The latter extracts the remainder 74% as quid pro quo payments.
While we have relied on some institutional features specific to SPACs, understanding how these features arise to alleviate information and agency frictions has implications for security issuance under asymmetric information in general and IPOs in particular. After all, the SPAC business combination process is also a case of security issuance through sponsors as intermediaries, with PIPE investors as buyers and the target company as the seller.
There are two main takeaways for IPOs. First, information production and quid pro quo can play important roles in IPO underpricing. In other words, they are not necessarily mutually exclusive. Second, non-premium PIPE investors help sponsors with weak deals, suggesting that, on average, they are value-enhancing for targets and issuers. This role is unlike the traditional IPO, where some investment bank clients get favorable allocations of “hot” IPO shares in exchange for paying high trading fees to the bank.
The benefit of quid pro quo payment is a function of its positive effect on successful business combinations. However, to determine if quid pro quo payments enhance welfare, we have to show that the total welfare is higher when the sponsor is allowed to build and deplete relationships than if not. These are interesting questions for future research.
@@ Appendix A Robustness
As shown in Figure A4, although our sample starts in 2010, most variation comes from SPAC from 2020 to 2022. Although our sample includes the so-called boom and bust cycle for SPAC, we may worry about the effects of outliers in our analysis. One such concern is that not all 34 premium PIPE investors (in Table 3) produce value-relevant information and that our results are driven by either the large premium PIPE investors or that the small premium PIPE investors are, in fact, non-premium investors.
@@
Note: This table presents estimates after excluding outliers in premium investors. Premium PIPE (at) investors are listed in Table 3. They include Fidelity Management & Research Company, BlackRock, Inc., Capital Research and Management Company, Alyeska Investment Group L.P., Morgan Stanley, and Millennium Management, LLC. Small premium investors are those that invested less than $0.4 billion and participated in at most 50 SPAC deals. They include Ghisallo Capital Management LLC, Jane Street Group, LLC, Park West Asset Management LLC, Kepos Capital LP, Linden Advisors LP, Schonfeld Strategic Advisors LLC, and BlueCrest Capital Management Ltd. Columns (1) and (2) have the same specification as Table 4, column (5). Columns (3) and (4) have the same specification as Table 5, column (5). Columns (5) and (6) have the same specification as Table 9, column (3). Columns (7) and (8) have the same specification as Table 11, column (5). , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
The results are presented in Table A1. In particular, the estimates from excluding large premium investors are in columns (1), (3), (5), and (7) of Table A1, and the estimates from excluding small premium investors are in the remaining columns. Comparing these estimates with those in columns (5), (5), (3), and (5) of Tables 4, 5, 9, and 11, respectively, we see that our main findings are neither due only to the large premium investors nor the small premium investors.

### v5

@@ 5.3.2 Allocations and Past Relationship
Next, we test the Hypothesis 3-3. A sponsor builds a relationship with non-premium investors by providing MLOT from good deals, so if and when the new target firm is weak, it expects the investors with high past MLOT, i.e., the relationship investors, to return the favor by investing in the current deal. To this end, we evaluate if a sponsor is more likely to use non-premium investors if their MLOT from the sponsor is high and the target is weak.
@@
Note: Columns (1) to (4) of this table show the MLE of participation (5), and column (5) shows the estimates of the allocation (6). Each observation is at (SPAC, sponsor, non-premium PIPE investor) level. For each SPAC, non-premium investors include all relationship investors who participate in at least one SPAC deal initialized by the sponsor during the sample period. For the first four columns, the dependent variable, , equals one if the non-premium investor participates in the focal SPAC deal. The dependent variable for the last column, , is the investor’s allocation as a fraction of all relationship investors’ allocations. The key independent variables include the non-premium investor’s MLOT earned with this sponsor in the past in log form, , and a measure of the (reciprocal) of the deal quality, Cold. For the allocation regression, we also include the estimated probability of participation using estimates from column (4) as an additional variable. All continuous variables at winsorized at percentile. Robust standard errors are clustered at the sponsor-PIPE pair level and reported in parentheses. For column (5), Bootstrap standard errors (500 replications) are reported in the square brackets. , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
One possible concern with the estimation is that it ignores the sponsor’s outside option, which may be correlated with the unobserved error. A better outside option lowers the probability of issuing securities to a non-premium investor and vice versa. Ignoring outside options could introduce attenuation bias if cold deals have profitable outside options. To alleviate this concern, we use the highest MLOT among all premium PIPE investors not included in the current deal as an additional control variable to capture the strength of the outside option. We also include the log of IPO proceeds, days remaining, and the number of premium investors as additional covariates.
@@
Next, we examine if, conditional on participation, the sponsor allocates a higher portion of securities to non-premium investors with a stronger relationship when a deal is cold and vice versa. To this end, we determine the allocation of SPAC ’s securities from sponsor to PIPE investor among all relationship investors , , as the ratio of ’s shares to the total shares allocated to . We set the share to zero if a non-premium investor does not participate in a deal. The mean allocation is , and the standard deviation is . Furthermore, the allocation is skewed towards zero, with the median of and the -percentile at . There are 1,792 non-premium investors, and the allocation has a mass at zero, so we estimate the following Tobit model:
@@ 5.3.3 Relationship Capital and Liquidation Risk
Our estimates so far suggest that the relationship investors, who are also non-premium investors, support sponsors by participating in weak deals. These relationship investors received larger MLOTs from the same sponsor in the previous deals than other non-premium investors. It stands to reason that a sponsor builds these relationships with an eye toward weak deals in the future–as insurance. If so, a sponsor with more “relationship capital” can afford to take more risk by working with weak targets.
@@
Next, we test the Hypothesis 3-4 by estimating the liquidation probability as a function of MLOT earned by non-premium investors in previous deals initiated by the same sponsor using a linear probability model.
@@
In our sample, there are 191 liquidations and 473 successful mergers. We verify if the sponsor’s “relationship payments” to non-premium investors in the SPAC lowers the probability of liquidation. To this end, we use a linear probability model in which the dependent variable is a binary variable equal to one if the SPAC is liquidated and zero otherwise. The explanatory variables include past MLOTs the non-premium investors have made from the current sponsor and several other variables.
@@
The results from this estimation exercise are shown in Table 11. Column (1) indicates SPACs initiated by sponsors with stronger relationships with non-premium investors are less likely to liquidate. We progressively introduce additional control variables from columns (2) to (6) as shown in Table 11. Column (2) includes the SPAC’s IPO month fixed effect to account for time differences. Column (3) incorporates fixed effects for the SPAC’s targeted sector at its IPO to capture systematic sectoral differences that may affect liquidation risk. Column (4) includes additional jurisdiction fixed effects based on the SPAC’s registration to capture legal system differences that may influence liquidation rules. Column (5) includes the sponsor fixed effect estimated from Equation (2) as a proxy for sponsor quality. The estimates show that SPAC’s liquidation risk decreases if its sponsor has a higher relationship with non-premium investors, and this effect is stable across different specifications.
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Past Relationship (Non-premium) | -0.064** | -0.074** | -0.077** | -0.079*** | -0.081*** | -0.065** |
| (0.029) | (0.030) | (0.030) | (0.030) | (0.031) | (0.030) | |
| Past Relationship (Premium) | 0.057* | 0.049 | 0.054* | 0.059* | 0.060* | 0.054* |
| (0.029) | (0.033) | (0.032) | (0.033) | (0.033) | (0.033) | |
| SPAC Return | -0.0001* | |||||
| {Premium Participation} | -0.372*** | |||||
| Controls | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| IPO Month FEs | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
| Target Sector FEs | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ |
| Jurisdiction FEs | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
| Sponsor Quality | ✗ | ✗ | ✗ | ✗ | ✓ | ✓ |
| Adj. | 0.048 | 0.121 | 0.126 | 0.126 | 0.125 | 0.254 |
| Obs. | 806 | 781 | 758 | 757 | 757 | 757 |
Note: This table shows that SPACs initiated by sponsors with stronger relationships with non-premium investors are less likely to liquidate. All regressions use linear probability models. The dependent variable is a dummy that equals one if the SPAC liquidates and zero otherwise. The main independent variables are the total MLOT earned by premium and non-premium investors in previous SPACs initiated by the same sponsor. SPAC Return is the return of holding the SPAC’s share from the offering date to one day before the merger completion date or the liquidation date. {Premium Participation} is a dummy variable that equals one if any premium investors participate in the focal deal. Controls include the (log) IPO proceeds, the number of past successful deals, a dummy variable for the presence of PIPE (pre) investors, and the number of active SPACs. Sponsor quality is measured by the sponsor’s fixed effect estimated from (2). All continuous variables are winsorized at the and percentiles. Robust standard errors, clustered at the sponsor level, are reported in parentheses. , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
A possible threat to the identification is that the set of non-premium investors could be endogenous if it is correlated with the unobserved quality of the target. If so, the estimated negative effect could be due to the positive selection between target and sponsor in that high-quality targets work with high-quality sponsors, who likely have completed more SPAC deals in the past and, therefore, have higher relationships with non-premium investors. So, the liquidation risk is lower because of the quality of the target. We include additional controls correlated with the target’s quality to address this issue. In particular, column (6) controls for the market return before the SPAC’s completion or liquidation and includes a dummy variable equal to one if a premium investor participates in the deal. Even with these new controls, we find that the estimated effect is negative and statistically significant at 5% confidence level.151515 Although not reported, we verify that our finding is robust to using the competing-risk model to account for right censoring because 16% of the announced deals in our sample are still searching for a target. In particular, the estimate suggests that doubling the sponsor’s relationship payments to non-premium investors decreases liquidation probability by 6.5%, which is 28.2% of the sample average liquidation rate, and this estimated effect is statistically significant at the 5% confidence level.161616On March 30, 2022, the SEC proposed new rules for SPAC mergers that made mergers more onerous and, consequently, relationships with non-premium investors more important. Indeed, estimating the specification in Table 11, column (6), separately for those SPACs that announced target before the law change and for those after, we find that the coefficient increases from to but are statistically indistinguishable.
@@
In summary, we find evidence consistent with these conditions: (i) premium PIPE investors generate value-relevant information, (ii) non-premium PIPE investors are paid agency fees, and (iii) the non-premium investors help close weaker deals in the future. In Appendix A, we assess the robustness of our results with respect to removing outlier investors, placebo treatments, and the definition of premium PIPE investors.
@@ 6 Conclusion
In this paper, we consider the relationship between intermediaries and investors in the securities issuance market. Our analysis distinguishes between premium and non-premium investors, revealing distinct roles for each investor type. Premium investors act in a certification or agency manner, whereas non-premium investors engage in relationships with intermediaries as Ljungqvist (2007) hypothesized. Specifically, when non-premium investors are allocated shares of hot deals, they are expected to reciprocate by supporting future lukewarm deals. These relationships serve as insurance for both the issuer and the intermediary.
We examine this insurance hypothesis in the context of SPACs, where the sponsor serves as an intermediary between the selling firm and PIPE investors. We focus on this setting because of its data advantages over traditional IPOs. With SPACs, we can directly observe stock price responses to public announcements, including the identity of the target and the participation decisions of PIPE investors. Additionally, we can directly observe the returns earned by each PIPE in each deal, allowing us to form a comprehensive panel data set of sponsor-PIPE relationships. Such sponsor-PIPE relationships data are unavailable in traditional IPOs, because there the initial allocations are proprietary information and investors often sell shares before filing a 13F, if at all.
Our findings indicate that premium investors are associated with higher announcement period returns, less redemption, and a higher likelihood of successful SPAC mergers. Conversely, non-premium investors exhibit mean-reverting returns, indicative of relationship capital buildup and drawdown. To establish the channel through which relationships are drawn down, we examine the participation and allocation of PIPE investors during cold deals, and show that these measures rise with past high returns. Independently of deal quality, sponsor-investor relationships significantly reduce the liquidation probability of current deals, demonstrating an insurance benefit accruing to intermediaries and issuers.
@@ A.1 Possible Outliers
As shown in Figure A2, although our sample starts in 2010, most of the variations come from SPAC from 2020 to 2022. Although our sample includes the so-called boom and bust cycle for SPAC, we may worry about the effects of outliers in our analysis. One such concern is that not all 34 premium PIPE investors (in Table 3) produce value-relevant information and that our results are driven by either the large premium PIPE investors or that the small premium PIPE investors are, in fact, non-premium investors.
@@
Note: This table presents estimates after excluding outliers in premium investors. Premium PIPE (at) investors are listed in Table 3. They include Fidelity Management & Research Company, BlackRock, Inc., Capital Research and Management Company, Alyeska Investment Group L.P., Morgan Stanley, and Millennium Management, LLC. Small premium investors are those that invested less than $0.4 billion and participated in at most 50 SPAC deals. They include Ghisallo Capital Management LLC, Jane Street Group, LLC, Park West Asset Management LLC, Kepos Capital LP, Linden Advisors LP, Schonfeld Strategic Advisors LLC, and BlueCrest Capital Management Ltd. Columns (1) and (2) have the same specification as Table 4, column (5). Columns (3) and (4) have the same specification as Table 5, column (5). Columns (5) and (6) have the same specification as Table 9, column (3). Columns (7) and (8) have the same specification as Table 11, column (6). , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
The results are presented in Table A1. In particular, the estimates from excluding large premium investors are in columns (1), (3), (5), and (7) of Table A1, and the estimates from excluding small premium investors are in the remaining columns. Comparing these estimates with those in columns (5), (5), (3), and (6) of Tables 4, 5, 9, and 11, respectively, we see that our main findings are neither due only to the large or small premium investors.